Fermi National Laboratory


Accelerator Update for November 18, 2005 - November 21, 2005

Friday November 18

The day shift began with Operations monitoring store 4506 and a small Antiproton Source (Pbar) stack, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, with the Main Injector (MI) down due to LCW leak repairs, with MiniBooNE taking beam, and with NuMI and the Meson MCenter and MTest experiments all down due to the MI work.

Operations reported at 8:42 AM that the MI work was completed and that the accelerator was being turned back on.

Pbar resumed stacking at 9:33 AM. NuMI also resumed taking beam.

At 9:33 AM, Operations reported that MiniBooNE was off due to a power supply problem (HV860). Operations contacted experts.

At 9:47 AM, a Linac Klystron RF station (KRF5) tripped off on a LCW flow indication for the first of four times in fifteen minutes.

At 10:23 AM, Operations reported that HV865 was also keeping MiniBooNE off. Experts continued their investigations.

Operations prepared the Recycler to stash antiprotons at 11:18 AM. The stash ended and Pbar resumed stacking at 12:26 PM. MTest and MCenter resumed taking beam at 12:30 PM.

A power supply expert reported at 1:22 PM that he had found and repaired a loose connection on HV865.

At 2:51 PM, Operations reported a P1 transfer line vacuum burst. It took about two minutes for the vacuum to recover.

At 3:51 PM, KRF5 tripped off on a flow trip. Operations had it back on line within five minutes, but then it tripped off again. Operators investigated.

The Recycler began preparing to stash at 5:57 PM. The stash ended and Pbar resumed stacking at 6:49 PM.

At 9:01 PM, a Booster RF station (BRF16) tripped off. It reset okay.

At 9:29 PM, BRF8 tripped off. It reset okay.

Operations terminated store 4506 at 10:04 PM. Tevatron (TeV) experts began a study period.

At 10:36 PM, BRF16 tripped off for the first of three times in about twenty minutes. It reset each time.

At 10:59 PM, KRF5 tripped off. Operators had it back on line within two minutes.

Saturday November 19

The midnight shift began with TeV experts conducting studies, with Operations monitoring the stash, with Pbar stacking, and with MiniBooNE, NuMI, MTest, and MCenter taking beam.

The TeV studies ended at 12:48 AM.

Operations put the TeV into shot setup at 12:57 AM.

At 1:19 AM, MiniBooNE tripped off due to HV865. Everything reset.

Operations established store 4508 at 3:11 AM with an initial luminosity of 126.4E30.

Operations received a sump alarm for MI-40. Operators prepared MI at 2:48 AM for an access. Operators reported that the area was dry. The high float alarm wasn’t working properly. The Duty Electrician estimated that it would take a couple of hours to troubleshoot. Operations decided to put off the repair to another time and resumed normal operations by 5:15 AM.

The Recycler began preparing to stash at 5:25 AM. The stash ended. Pbar, MTest, and MCenter resumed normal operations.

At 8:03 AM, HV865 tripped off. It reset okay.

At 9:29 AM, the TeV suffered a sector D, E, and F quench. Store 4508 was lost. Operators determined that the cause was due to a sector B11 separator spark.

The Recycler began preparing to stash at 10:17 AM. One transfer was lost due to a P150 transfer line vacuum burst.

Operations prepared the D0 and CDF collision halls for access at 10:49 AM.

The stash ended. Pbar, MTest, and MCenter resumed normal operations at 11:08 AM. Both experiments were out of access by 1:29 PM.

Operations put the TeV through a dry squeeze at 1:37 PM.

TeV experts began a study period at 2:14 PM.

The Recycler began preparing to stash at 3:25 PM.

At 4:01 PM, Operations reported that MCenter had accessed their hall during the stash. Something happened to the interlock system during the access. Operations called in interlock technicians who made repairs. MCenter resumed taking beam about 2.5 hours later.

The stash ended. Pbar and MTest resumed normal operations by 4:15 PM.

The Recycler began preparing to stash at 8:31 PM. The stash ended. Pbar, MTest, and MCenter resumed normal operations by 8:57 PM.

Operations put the TeV through a dry squeeze at 9:13 PM.

At 9:23 PM, Operations reported a TeV sector A0 vacuum valve that wouldn’t open. Operators investigated and then called in an expert. The expert replaced a faulty valve control card.

At 10:45 PM, Operations reported that there was a B11 separator spark during the dry squeeze.

Operation put the TeV into a wet squeeze at 11:40 PM.

Sunday November 20

The midnight shift began with the TeV in a wet squeeze, Operations monitoring the stash, with Pbar stacking, and with MiniBooNE, NuMI, MTest, and MCenter taking beam.

Operations put the TeV into shot setup around 1 AM.

Operations established store 4515 at 2:58 AM with an initial luminosity of 124E30.

The Recycler began preparing to stash at 2:59 AM. The stash ended. Pbar, MTest, and MCenter resumed normal operations by 3:44 AM.

The Recycler began preparing to stash at 8:20 AM. The stash ended. Pbar, MTest, and MCenter resumed normal operations by 8:55 AM.

The Recycler began preparing to stash at 1:32 PM. The stash ended. Pbar, MTest, and MCenter resumed normal operations by 2:10 PM.

Operations prepared Pbar for a reverse proton study at 4:03 PM. During this time an expert tuned the Booster beam.

At 5:01 PM, the P150 transfer line suffered a vacuum burst.

At 9:07 PM, LRF2 tripped off. It took a number of resets, but operators managed to get it back on line.

The reverse proton study ended and Pbar resumed stacking at 9:20 PM.

At 11:38 PM, LRF2 tripped off twice in two minutes. It reset okay.

At 11:53 PM, LRF2 tripped off several times and then wouldn’t reset. Operators investigated.

Monday November 21

The midnight shift began with Operations monitoring store 4515 and the stash, with all beam off due to LRF2.

Operators had LRF2 reset at 12:03 AM. Beam returned to all experiments.

The Recycler began preparing to stash at 3:10 AM. The stash ended. Pbar, MTest, and MCenter resumed normal operations by 4:26 AM.

At 4:39 AM, Operations reported that their alarm program stopped working. After trying several things, Operations called an expert.

Operations terminated store 4515 at 5:12 AM.

Operations put the TeV into shot setup at 5:30 AM.

At 6:31 AM, Pbar stopped stacking. Operators investigated and discovered that a power supply (H717) LCW leak in the F27 service building had drained enough water to cause the LCW pumps to trip off and take down power supplies for the AP2 & 3 lines. Operators called in power supply experts and contacted a water expert.

The Plans for Monday

The plans for today are to stash, stack, and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 11/21/2005   email Fermilab